OpenStudy (anonymous):

A 89-year-old person is 7.7964 × 105 hours old. Write this number in standard form.

OpenStudy (anonymous):

7.7964x5x10^2 =3.8982x10^3

OpenStudy (anonymous):

105 = 100 + 5 => 100 = 10^2. How did you get 7.7964x5x10^2

OpenStudy (anonymous):

5x10^2 = 500 isn't it?

OpenStudy (anonymous):

lol